OpenStudy (anonymous):

What is the equation of the line pretriceing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

from point (-6,1),,, the slope perpendicular to y = –3x + 1 is +3 or m=1/3 using y=mx+b 1=(1/3)(-6)+b 1=-2+b 1+2=b b=3 now y=mx+b y=(1/3)x+3 ans